The lateral area of a right cylinder whose height is 3 times the length of the radius is 486 pi. What is the height of the cylinder?
Answer by cleomenius(959)

The way I read this problem, you would calculate this as you would a rectangle,
A = B * H
3x^2 * 2x = 6x^2 = 486
x^2 = 81
x = 9
18 = width, 27 = height.
Cleomenius.